A side of an equilateral triangle is 40 m long. What is the area of the triangle?
Answer by nerdybill(7384)

A side of an equilateral triangle is 40 m long. What is the area of the triangle?
.
area is (1/2)*base*height
the only thing we need is the height -- find it by applying Pythagorean theorem:
h^2 + 10^2 = 20^2
h^2 + 100 = 400
h^2 = 300
h = sqrt(300)
h = sqrt(10*10*3)
h = 10*sqrt(3)
.
area = (1/2)(20)(10*sqrt(3))
area = (10)(10*sqrt(3))
area = 100*sqrt(3)
area = 173.21 sq meters